# Federated Unlearning: a Perspective of Stability and Fairness

### v1

@@ 7.1 Experiment Settings
In our experiment, we utilize the MNIST dataset LeCun et al. non-IID distributed across ten clients, each holding four distinct classes (the data distribution is detailed in Appendix G). We employ LeNet-5 architecture LeCun et al. (1998), a classic non-convex neural network model, to evaluate FU’s consequences. To straightforwardly assess the FU evaluations metrics ($V,S,Q$), we focus on accuracy, e.g., $V=-Acc_{-\mathcal{J}}(\boldsymbol{{w}}^{u})+Acc_{-\mathcal{J}}(\boldsymbol{w}^{r*})$ (in percentage). A smaller value of these metrics indicates better effectiveness, stability, or fairness achieved by our FU mechanism. We further conduct additional experiments focusing on data heterogeneity and employ different datasets, as detailed in Appendix G. The overall experimental evaluation confirms our FU mechanisms in balancing the trade-offs, aligning with the theoretical insights in Section 5.
@@ 7.2 FU for Balancing Stability
Unlearning Convergence and Handling Stability. The heterogeneity between remaining and unlearned clients leads to instability after unlearning (discussed in Lemma 5.8), as indicated by the reduced global performance after retraining in Figure 4. Additionally, Figure 4 showcases the convergence of our FU mechanism with different stability penalties $\lambda$ in the context of global performance. Specifically, with a stability penalty $\lambda=1$, unlearning shows better stability than exact retraining. Increasing $\lambda$ to 5 further improves the stability of FU.
Balancing Verification and Stability: As shown in Figure 5, increasing $\lambda$ lowers $V+S$, improving the balance between verification and stability in FU. Although a higher $\lambda$ reduces FU effectiveness (increasing $V$), our FU mechanism allows a better trade-off within a certain tolerance level for $V$. Additionally, it demonstrates time efficiency compared to retraining in Table 1, making it practical in real-world FU scenarios.
Figure 5: Trade-offs with Stability Penalty $\lambda$: Increasing $\lambda$ improves the balance between FU verification and stability, reflected by reduced $V+S$.
Table 1: Unlearning Efficiency Compared to Retraining.
(Computing Resources: 2 Intel Xeon Gold 5217 CPUs, 384GB RAM, and 8 Nvidia GeForce RTX-2080Ti GPUs)
@@ 7.3 FU for Balancing Fairness
Considering a higher data heterogeneity between unlearned and remaining clients, unlearning clients [2, 10] leads to a maximum performance drop of 1.5% ($\epsilon=1.5$) for client 6. We employ the FU mechanism in Section 6.2 with a fairness constraint $\epsilon=1$ and $\Lambda=1.3$ ensuring no client’s performance deviates beyond $\epsilon$. This setting yields the FU verification metric $V=0.57$. We further tighten $\epsilon$ to $1$, and observe $V$ reduces to $0.11$. This enhancing FU effectiveness and fairness is consistent with insights on data heterogeneity in Lemma 6.8 and Theorem 5.11.
@@ 8 Conclusion
$\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{t+1})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{t})\geq\frac{\eta}{2}\underbrace{\left\|\boldsymbol{G}_{\mathcal{S}}^{t}-\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}}_{A_{1}}\underbrace{-\frac{\eta}{2}\|\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\|^{2}-\frac{\eta}{2}\left\|\boldsymbol{G}_{\mathcal{S}}^{t}\right\|^{2}+\frac{\mu\eta^{2}}{2}\mathbb{E}\|\boldsymbol{g}_{\mathcal{S}}^{t}\|^{2}}_{A_{2}}$ (11)
@@
[…]
@@
Taking $\eta^{2}=\frac{2}{\mu MT^{2}}$, we have $B_{1}=\frac{2(\beta_{\mathcal{S}}^{2}-1)}{T^{2}}$.
@@
Case 1: If $B_{1}\leq-1$ (since $T^{2}\geq 1$, Case 1 only holds when $T=1$ and $\beta_{\mathcal{S}}^{2}\leq\frac{1}{2}$), we get:
$\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{T})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\geq\left(1+2(\beta_{\mathcal{S}}^{2}-1)\right)\left(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\right)+\frac{1}{M}\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$ (17)
Case 2: If $B_{1}\geq-1$, then $(B_{1}+1)^{T}\geq 1+TB_{1}$, which leads to the inequality:
$\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{T})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\geq\left(1+2\frac{\beta_{\mathcal{S}}^{2}-1}{T}\right)\left(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\right)+\rho\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$ (18)
@@
$~{}\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{T})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})$ $\geq\left(1+2\frac{\beta_{\mathcal{S}}^{2}-1}{T}\right)\left(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\right)+\frac{1}{2LT}\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$ (19)
@@
With $\boldsymbol{{w}}^{u}-\boldsymbol{\boldsymbol{w}}^{o}=-\eta\sum_{r=1}^{T}\mathbf{g}_{r}^{(\mathcal{S})}$, and $\mathbf{g}_{r}^{(\mathcal{S})}$ being the aggregated stochastic gradient from the subset of remaining clients $\mathcal{S}\subseteq\mathcal{N}\setminus\mathcal{J}$, let $\bar{\mathbf{g}}^{(\mathcal{S})}=\frac{1}{T}\sum_{r=1}^{T}\mathbf{g}_{r}^{(\mathcal{S})}$, we have $\boldsymbol{{w}}^{u}-\boldsymbol{\boldsymbol{w}}^{o}=-\eta T\bar{\mathbf{g}}^{(\mathcal{S})}$.
@@
$=\frac{(1-P_{\mathcal{J}})\eta T}{2}\mathbb{E}\left[\|\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\bar{\mathbf{g}}^{(\mathcal{S})}\|^{2}\right]+\frac{P_{\mathcal{J}}\eta T}{2}\mathbb{E}\left[\|\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\bar{\mathbf{g}}^{(\mathcal{S})}\|^{2}\right]$
@@
$\quad-\frac{\eta T}{2}\mathbb{E}\left[\|\bar{\mathbf{g}}^{(\mathcal{S})}\|^{2}\right]+\frac{\mu\eta^{2}T^{2}}{2}\mathbb{E}\left[\|\bar{\mathbf{g}}^{(\mathcal{S})}\|^{2}\right],$ (21)
@@
$A_{1}$ $\geq\frac{P_{\mathcal{J}}\eta T}{2}\|\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})\|^{2}+\frac{\eta^{2}T^{2}-\mu T}{2}(\|\bar{\mathbf{G}}^{(\mathcal{S})}\|^{2}+\sigma^{2})+\frac{\eta T}{2}G^{2}$
@@ Appendix E Balancing Stability Unlearning Algorithm Analysis
Lemma 1 of Li et al. (2020b) Let 5.1 holds and $\eta_{l}\leq\frac{1}{4L}$, the remaining clients’ contribution to unlearning for every round $t$ gives:
@@
where $R^{t}$ can be bounded by Lemma E.1.
@@
where $\Omega=(\beta^{\prime 2}+1)$, $\beta^{\prime 2}$ indicates the data heterogeneity between remaining and unlearned clients.
@@
By the $L$-smoothness of $H(\cdot)$,
$\mathbb{E}\left[H\left({\boldsymbol{\bar{w}}}^{(t,0)}\right)\right]$ $-H^{*}$
$=\mathbb{E}\left[H\left({\boldsymbol{\bar{w}}}^{(t,0)}\right)\right]-H(\boldsymbol{w}^{r*})+H(\boldsymbol{w}^{r*})-H^{*}$
$\leq\frac{L}{2}\Delta_{t}+\frac{L}{2}\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{*}\right\|^{2}$
$\leq\frac{L}{2}\left(\frac{v}{\gamma+t}+\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{*}\right\|^{2}\right)$
@@
$\frac{1}{2}\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{t+1})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{t})$ $\leq A_{1}=\frac{\eta_{l}}{2}\mathbb{E}\left[\left\|\boldsymbol{g}_{\mathcal{S}}^{t}-\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}\right]-\frac{\eta_{l}}{2}{\left\|\boldsymbol{G}_{\mathcal{S}}^{t}\right\|^{2}}-\frac{\eta_{l}}{2}\left\|\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}+\frac{L\eta_{l}^{2}}{2}\left\|\boldsymbol{g}_{\mathcal{S}}^{t}\right\|^{2}$

### v4

@@ 7.1 Experiment Settings
Figure 4: Non-IID Class Distribution for MNIST and CIFAR-10: Each client has four classes.
In our experiment, our implementation utilizes the open-source FL framework Plato [25] for reproducibility. We utilize the MNIST dataset [24] and the CIFAR-10 dataset [21] non-IID distributed across ten clients, each holding four distinct classes (the data distribution is detailed in Figure 4).
We employ LeNet-5 architecture [23] and ResNet18 architecture [16], classic non-convex neural network models, to evaluate FU’s side effects. To straightforwardly assess the FU evaluations metrics ($V,S,Q$), we focus on accuracy, e.g., $V=-Acc_{-\mathcal{J}}(\boldsymbol{{w}}^{u})+Acc_{-\mathcal{J}}(\boldsymbol{w}^{r*})$ (in percentage). A smaller value of these metrics indicates better effectiveness, stability, or fairness achieved by our FU mechanism. The overall experimental evaluation confirms our FU mechanisms in balancing the trade-offs, aligning with the theoretical insights in Section 5.
@@ 7.2 FU for Balancing Stability
(a) Unlearning Convergence and Handling Stability: Increasing stability penalty $\lambda$ enhances global performance of FU. (MNIST + LeNet-5)
(b) Trade-offs with Stability Penalty $\lambda$: Increasing $\lambda$ improves the balance between FU verification and stability, reflected by reduced $V+S$.
Figure 5: Comparison of stability-related processes.
Table 1: Balancing Stability in FU (ResNet18, CIFAR10)
Global Performance (%) Unlearning Performance (%) V (%) S (%) V + S (%)
Original FL 63.6 - - - -
Retrain 61.18 63.99 0 2.42 2.42
Unlearning ($\lambda=1$) 62.10 63.57 0.42 1.42 1.84
Unlearning Convergence and Handling Stability. The heterogeneity between remaining and unlearned clients leads to instability after unlearning (discussed in Lemma 5.8), as indicated by the reduced global performance after retraining in Figure 5(a). Additionally, Figure 5(a) showcases the convergence of our FU mechanism with different stability penalties $\lambda$ in the context of global performance. Specifically, with a stability penalty $\lambda=1$, unlearning shows better stability than exact retraining (Figure 5(a), Table 1). Increasing $\lambda$ to 5 further improves the stability of FU.
Balancing Verification and Stability: As shown in Figure 5(b), increasing $\lambda$ lowers $V+S$, improving the balance between verification and stability in FU.
Although a higher $\lambda$ reduces FU effectiveness (increasing $V$), our FU mechanism allows a better trade-off within a certain tolerance level for $V$. Additionally, it demonstrates time efficiency compared to retraining in Table 2, making it practical in real-world FU scenarios.
Table 2: Unlearning Efficiency Compared to Retraining.(Computing Resources: 2 Intel Xeon Gold 5217 CPUs, 384GB RAM, and 8 Nvidia GeForce RTX-2080Ti GPUs)
@@ 7.3 FU for Balancing Fairness
Figure 6: Unlearning Impact on Fairness (CIFAR10 + ResNet18): ensuring no client’s performance deviates beyond $20\%$; fairness metric ($Q$) of our FU mechanism ($Q_{\text{our}}=6.91$) is lower than that of retrain $Q_{\text{retrain}}=10.23$
This section explores the fairness in FU under varying levels of heterogeneity among remaining clients, and data between remaining and unlearned clients is also heterogeneous [^5] According to Lemma 6.8, given homogeneous two groups, the fairness primarily depends on data heterogeneity among remaining clients, as we aim to investigate the impact of unlearned clients, we choose heterogeneous two groups setting.. Figure 6 illustrates the impact of unlearning on fairness among remaining clients, with a fairness parameter $\Lambda=1,\epsilon=20$. In this scenario, clients [1, 3] experienced significant utility loss due to their similar data distribution with unlearned clients [8, 9] (as shown in Figure 4). However, our FU mechanism in Section 6.2 achieves a lower fairness metric ($Q_{\text{our}}=6.91$) compared to the retraining approach ($Q_{\text{retrain}}=10.23$), indicating more equitable utility changes among remaining clients. The verification metric in this case is $V=0.42$, demonstrating that our mechanism enhances fairness even in more complex and heterogeneous environments. This enhancing FU effectiveness and fairness is consistent with insights on data heterogeneity in Lemma 6.8 and Theorem 5.11.
7.4 Heterogeneity between remaining and unlearned clients
Table 3 demonstrates a correlation between the data heterogeneity level and the system’s stability after unlearning. A higher data heterogeneity leads to increased instability after unlearning, as indicated by the higher $S_{\text{retrain}}$ values. Notably, applying our FU mechanism with a stability trade-off parameter ($\lambda=1$) results in improved stability, as shown by the reduced $S_{\lambda=1}$ values, while the difference in FU verification remains negligible. This underscores our proposed approach could balance the trade-off between unlearning verification and global stability under varying degrees of data heterogeneity.
Label Distribution $S_{\text{retrain}}$ $S_{\lambda=1}$ $V_{\lambda=1}$
0.1 8.24 7.69 (-0.55) 0.0
0.4 3.45 2.05 (-1.4) 0.0
0.7 0.26 0.23 (-0.03) 0.06
@@ 8 Conclusion
$\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{t+1})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{t})\geq\frac{\eta}{2}\underbrace{\left\|\boldsymbol{G}_{\mathcal{S}}^{t}-\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}}_{A_{1}}$
$\underbrace{-\frac{\eta}{2}\|\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\|^{2}-\frac{\eta}{2}\left\|\boldsymbol{G}_{\mathcal{S}}^{t}\right\|^{2}+\frac{\mu\eta^{2}}{2}\mathbb{E}\|\boldsymbol{g}_{\mathcal{S}}^{t}\|^{2}}_{A_{2}}$ (9)
@@
$A_{2}=\frac{\mu\eta^{2}}{2}\underbrace{\mathbb{E}\left\|\boldsymbol{G}_{\mathcal{S}}^{t}-\boldsymbol{g}_{\mathcal{S}}^{t}\right\|^{2}}_{=\sigma_{t}^{2}}+\frac{\eta}{2}\left(\eta\mu-1\right)\left\|\boldsymbol{G}_{\mathcal{S}}^{t}\right\|^{2}-\frac{\eta}{2}\left\|\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}$ (11)
@@
Taking $\eta^{2}=\frac{2}{\mu MT^{2}}$, we have $B_{1}=\frac{\beta_{\mathcal{S}}^{2}-1}{T^{2}}$.
@@
Case 1: If $B_{1}\leq-1$ (since $T^{2}\geq 1$, Case 1 only holds when $T=1$ and $\beta_{\mathcal{S}}^{2}=0$), we get:
$\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{T})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\geq\frac{1}{M}\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$ (15)
Case 2: If $B_{1}>-1$, then $(B_{1}+1)^{T}\geq 1+TB_{1}$, which leads to the inequality:
$\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{T})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\geq\left(1+\frac{\beta_{\mathcal{S}}^{2}-1}{T}\right)\left(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\right)+\rho\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$ (16)
@@
$~{}\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{T})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})$ $\geq\left(1+\frac{\beta_{\mathcal{S}}^{2}-1}{T}\right)\left(F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-F_{-\mathcal{J}}(\boldsymbol{w}^{r*})\right)+\frac{1}{2LT}\left(\bar{\sigma^{2}}+\bar{\zeta^{2}}\right),$ (17)
@@
With $\boldsymbol{{w}}^{u}-\boldsymbol{\boldsymbol{w}}^{o}=-\eta\sum_{r=1}^{T}\mathbf{g}_{\mathcal{S}}^{r}$, and $\mathbf{g}_{\mathcal{S}}^{r}$ being the aggregated stochastic gradient from the subset of remaining clients $\mathcal{S}\subseteq\mathcal{N}\setminus\mathcal{J}$, let $\bar{\mathbf{g}}_{\mathcal{S}}=\frac{1}{T}\sum_{r=1}^{T}\mathbf{g}_{\mathcal{S}}^{r}$, we have $\boldsymbol{{w}}^{u}-\boldsymbol{\boldsymbol{w}}^{o}=-\eta T\bar{\mathbf{g}}_{\mathcal{S}}$.
@@
$=\frac{(1-P_{\mathcal{J}})\eta T}{2}\mathbb{E}\left[\|\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\bar{\mathbf{g}}_{\mathcal{S}}\|^{2}\right]+\frac{P_{\mathcal{J}}\eta T}{2}\mathbb{E}\left[\|\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\bar{\mathbf{g}}_{\mathcal{S}}\|^{2}\right]$
@@
$\quad-\frac{\eta T}{2}\mathbb{E}\left[\|\bar{\mathbf{g}}_{\mathcal{S}}\|^{2}\right]+\frac{\mu\eta^{2}T^{2}}{2}\mathbb{E}\left[\|\bar{\mathbf{g}}_{\mathcal{S}}\|^{2}\right],$ (19)
@@
$A_{1}$ $\geq\frac{P_{\mathcal{J}}\eta T}{2}\|\nabla F_{-\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})-\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})\|^{2}+\frac{\eta^{2}T^{2}-\mu T}{2}(\|\bar{\mathbf{g}}_{\mathcal{S}}\|^{2}+\sigma^{2})+\frac{\eta T}{2}G^{2}$
@@ Appendix E Balancing Stability Unlearning Algorithm Analysis
(Lemma 1 of [29]) Let 5.1 holds and $\eta_{l}\leq\frac{1}{4L}$, the remaining clients’ contribution to unlearning for every round $t$ gives:
@@
where $R^{t}={\mathbb{E}\left[\left\|{\boldsymbol{\bar{w}}}^{(t,E)}-\boldsymbol{w}^{r*}\right\|^{2}\right]}$ can be bounded by Lemma E.1.
@@
where $\Omega=(\beta^{\prime 2}+1)$, and $\beta^{\prime 2}$ indicates the data heterogeneity between remaining and unlearned clients.
@@
By the $L$-smoothness of $H(\cdot)$ and AM-GM inequality and Cauchy–Schwarz inequality, we can derive:
$\mathbb{E}\left[H\left({\boldsymbol{\bar{w}}}^{(t,0)}\right)\right]$ $-H^{*}=\mathbb{E}\left[H\left({\boldsymbol{\bar{w}}}^{(t,0)}\right)\right]-H(\boldsymbol{w}^{r*})+H(\boldsymbol{w}^{r*})-H^{*}$
$\leq\left({\boldsymbol{\bar{w}}}^{(t,0)}-\boldsymbol{w}^{r*}\right)^{\top}\nabla H(\boldsymbol{w}^{r*})+\frac{L}{2}\Delta_{t}+\frac{L}{2}\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}$
$\leq\frac{L}{2}\Delta_{t}+\frac{1}{2L}\left\|H(\boldsymbol{w}^{r*})\right\|^{2}+\frac{L}{2}\Delta_{t}+\frac{L}{2}\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}$
$=\frac{1}{2L}\left\|P_{\mathcal{J}}\lambda\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})+L(\boldsymbol{w}^{r*}-\boldsymbol{\boldsymbol{w}}^{o})\right\|^{2}+L\Delta_{t}+\frac{L}{2}\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}$
$\leq\frac{1}{2L}\left\|P_{\mathcal{J}}\lambda\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})\right\|^{2}+\frac{L}{2}\left\|\boldsymbol{w}^{r*}-\boldsymbol{\boldsymbol{w}}^{o}\right\|^{2}+L\Delta_{t}+\frac{L}{2}\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}$
$\leq L\frac{v}{\gamma+t}+\frac{1}{2L}\left\|P_{\mathcal{J}}\lambda\nabla F_{\mathcal{J}}(\boldsymbol{\boldsymbol{w}}^{o})\right\|^{2}+\frac{L}{2}\left(\left\|\boldsymbol{w}^{r*}-\boldsymbol{\boldsymbol{w}}^{o}\right\|^{2}+\left\|\boldsymbol{w}^{r*}-\boldsymbol{w}^{H*}\right\|^{2}\right)$
@@
To prove Lemma 5.6, we begin to introduce and prove the following additional lemmas:
@@
Lemma E.4.
Assume 6.1 holds. Given a set of clients $\mathcal{S}$, it follows that
$\mathbb{E}\left\|\boldsymbol{G}_{\mathcal{S}}^{t}-\boldsymbol{g}_{\mathcal{S}}^{t}\right\|^{2}\leq\sum_{i\in\mathcal{S}}\alpha_{i}^{2}\sigma_{i,t}^{2},$
where $\boldsymbol{G}_{\mathcal{S}}^{t}=\mathbb{E}\left[\boldsymbol{g}_{\mathcal{S}}^{t}\right]$.
@@
$\frac{1}{2}\mathbb{E}\left[F_{-\mathcal{J}}(\boldsymbol{w}^{t+1})\right]-F_{-\mathcal{J}}(\boldsymbol{w}^{t})\leq A_{1}$
$=\frac{\eta_{l}}{2}\mathbb{E}\left[\left\|\boldsymbol{g}_{\mathcal{S}}^{t}-\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}\right]-\frac{\eta_{l}}{2}{\left\|\boldsymbol{G}_{\mathcal{S}}^{t}\right\|^{2}}-\frac{\eta_{l}}{2}\left\|\nabla F_{-\mathcal{J}}(\boldsymbol{w}^{t})\right\|^{2}+\frac{L\eta_{l}^{2}}{2}\left\|\boldsymbol{g}_{\mathcal{S}}^{t}\right\|^{2}$
@@
, where $\sigma^{2}=\sum_{i\in\mathcal{S}}\alpha_{i}^{2}\sigma_{i,t}^{2}$ by Lemma E.4.